OpenStudy (anonymous):

Find the solutions of the inequality. | x - 5 | ≥ 3

OpenStudy (anonymous):

Honestly, Why don't you at least attempt to solve it?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

its 8

OpenStudy (anonymous):

Do you mind posting it?

OpenStudy (anonymous):

:] wooh i feeel smart :3

OpenStudy (anonymous):

omg .. =..=

OpenStudy (anonymous):

its \[x \ge 8\]

OpenStudy (anonymous):

2<re(x)<8

OpenStudy (anonymous):

Now i feel smater :}

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

re(x)>8

OpenStudy (anonymous):

There are two possible solutions.

OpenStudy (anonymous):

\[| x - 5 | ≥ 3\] \[=> x-5 \ge 3 ..or.. x-5 \le -3\] NOw solve it yourself to get the answer :)